# Driving with DINO: Vision Foundation Features as a Unified Bridge for Sim-to-Real Generation in Autonomous Driving

### v1

@@ 4.5. Ablation Studies
Table 2. Ablation studies. $\times S$ means the upscaling spatial resolution $S$ times.
@@
Table 3. Comparison of mIoU across different methods and categories.
@@
[…]
5. Conclusion
@@
Figure 6. Extra Qualitative comparison with state-of-the-art methods. The red boxes highlight inconsistencies with respect to the CG input, while the green boxes point out low-fidelity textures.
Figure 7. Extra Qualitative comparison with state-of-the-art methods. The red boxes highlight inconsistencies with respect to the CG input, while the green boxes point out low-fidelity textures.
@@ Appendix A Implementation details
A.1. Training Details
We leverage the Nuplan dataset (caesar2021nuplan) to train DwD. To address the inconsistent quality of raw sequences, we implement an automated data curation pipeline utilizing Qwen3-VL (Qwen3-VL). This pipeline filters out samples exhibiting visual degradation—such as over-saturation, monochromatic artifacts, or other anomalies—and generates captions for the remaining data. Consequently, we compile a curated dataset of 6,000 videos, each containing 200 frames.To ensure a stable input space for DwD, we address the inherent sign ambiguity of PCA eigenvectors by establishing a consistent global feature space via Incremental PCA. Specifically, we adopt a sparse sampling strategy that extracts a single random frame from each video to compute a fixed global basis (comprising the mean and eigenvectors). This approach maximizes environmental diversity (e.g., weather conditions, time of day) while mitigating sequence-specific biases. All feature embeddings are projected onto this basis prior to training.During training, video sequences are resized to a resolution of $1280\times 704$ and randomly cropped into temporal chunks of 93 frames. We adopt a focused fine-tuning strategy: the parameters of the original denoising DiT are frozen, while only the copied conditional DiT is updated. Due to memory constraints (48GB VRAM), we employ Context Parallelism on a single node with 8 GPUs. The model is trained for 20k iterations with a total batch size of 1 and a learning rate of $5\times 10^{-5}$.
A.2. Data rendered from Simulator
@@
A.3. Spatial Alignment Module and Causal Temporal Aggregator
@@
Figure 8. The Spatial Alignment Module and Causal Temporal Aggregator.
@@ Appendix B Qualitative Comparisons across Ablated Hyperparameters
Figure 9. Qualitative Comparison across Minor Components Pruning Degree ($k$)
Figure 10. Qualitative Comparison across different Upscaling Factors $\times S$ in Spatial Resolution Enhancement. The red boxes highlight inconsistencies with respect to the CG input.
Figure 11. Qualitative comparison of autoregressive long video generation.
B.1. Analysis of Naive DINO Latent Injection
@@
B.2. Impact of Minor Components Pruning Degree ($k^{\prime}$)
@@
B.3. Effect of Spatial Upscaling Factor ($S$)
@@
Figure 12. Qualitative Comparison w. and w./o. Causal Temporal Aggregator
Figure 13. PCA visualization comparing feature maps upsampled via AnyUp ($\times 4$ and $\times 8$) versus naive input upscaling($\times 4$). The red boxes highlight semantic inconsistencies introduced by AnyUp, where high-frequency pixel information leads to false semantic boundaries (e.g., misclassifying the curb). The green boxes demonstrate the failure of AnyUp to recover fine-grained structural details, such as the boundaries of distant street lamps, despite the high upsampling factor.
B.4. Efficacy of the Causal Temporal Aggregator
@@ Appendix C Upsampling using Anyup
While feature upsampling methods such as AnyUp (wimmer2025anyup) appear to offer a computationally efficient alternative for obtaining high-resolution DINO feature maps, we empirically find that they consistently lead to unstable controllability in video generation, resulting in the suboptimal quantitative metrics and poor visual quality. To investigate this, we visualize the PCA components of feature maps upsampled via AnyUp versus naive input upscaling in Fig. 13.We observe that since AnyUp incorporates high-resolution RGB images as guidance, it inadvertently introduces unnecessary high-frequency pixel information, leading to feature quality degradation. As highlighted by the red boxes, the pixel-level guidance causes AnyUp to misinterpret texture differences (e.g., between the curb and the neighboring ground) as distinct semantic categories, whereas naive upscaling maintains a more coherent semantic representation. Furthermore, as shown in the green boxes, AnyUp fails to recover fine-grained structural details despite an $8\times$ upsampling factor; for instance, the boundaries between distant street lamps and the sky remain indistinct. These visual observations corroborate our experimental findings, confirming that the feature degradation inherent in AnyUp is the primary cause of the generation instability.

### v2

@@ 4.5 Ablation Studies
Table 2: Ablation studies. $\times S$ means the upscaling spatial resolution $S$ times.
@@
Table 3: Comparison of mIoU across different methods and categories.
@@
Road 0.9574 0.9591 0.9636 0.9475 0.9345 0.9605
Sidewalk 0.5464 0.5914 0.6332 0.5253 0.5010 0.5610
Building 0.7802 0.8193 0.8127 0.7725 0.5781 0.8087
Fence 0.0978 0.0767 0.0790 0.0789 0.0848 0.1187
Pole 0.2757 0.3665 0.3214 0.3548 0.2555 0.3435
Traffic Light 0.4354 0.6655 0.5501 0.5392 0.3174 0.4503
Traffic Sign 0.2014 0.3838 0.2862 0.2601 0.1505 0.2681
Vegetation 0.6907 0.6833 0.7210 0.6936 0.5773 0.7150
Terrain 0.3171 0.3561 0.4476 0.3284 0.3210 0.3245
Sky 0.8945 0.8912 0.9028 0.9115 0.7691 0.9111
Person 0.5380 0.6795 0.6729 0.6438 0.5828 0.6256
Car 0.7522 0.7973 0.7819 0.7828 0.7044 0.7756
Bus 0.0989 0.0223 0.0620 0.0259 0.0049 0.0296
Bicycle 0.0599 0.0242 0.0132 0.0294 0.0612 0.0666
mIoU (Avg) 0.4675 0.5015 0.5105 0.4843 0.4132 0.4967
5 Conclusion
@@ Appendix A Implementation details
A.1 Training Details
We leverage the Nuplan dataset [5] to train DwD. To address the inconsistent quality of raw sequences, we implement an automated data curation pipeline utilizing Qwen3-VL [3]. This pipeline filters out samples exhibiting visual degradation—such as over-saturation, monochromatic artifacts, or other anomalies—and generates captions for the remaining data. Consequently, we compile a curated dataset of 6,000 videos, each containing 200 frames.To ensure a stable input space for DwD, we address the inherent sign ambiguity of PCA eigenvectors by establishing a consistent global feature space via Incremental PCA. Specifically, we adopt a sparse sampling strategy that extracts a single random frame from each video to compute a fixed global basis (comprising the mean and eigenvectors). This approach maximizes environmental diversity (e.g., weather conditions, time of day) while mitigating sequence-specific biases. All feature embeddings are projected onto this basis prior to training.During training, video sequences are resized to a resolution of $1280\times 704$ and randomly cropped into temporal chunks of 93 frames. We adopt a focused fine-tuning strategy: the parameters of the original denoising DiT are frozen, while only the copied conditional DiT is updated. Due to memory constraints (48GB VRAM), we employ Context Parallelism on a single node with 8 GPUs. The model is trained for 20k iterations with a total batch size of 1 and a learning rate of $5\times 10^{-5}$.
A.2 Data rendered from Simulator
@@
A.3 Spatial Alignment Module and Causal Temporal Aggregator
@@
Figure 5: The Spatial Alignment Module and Causal Temporal Aggregator.
@@ Appendix B Qualitative Comparisons across Ablated Hyperparameters
Figure 6: Qualitative Comparison across Minor Components Pruning Degree ($k$)
Figure 7: Qualitative Comparison across different Upscaling Factors $\times S$ in Spatial Resolution Enhancement. The red boxes highlight inconsistencies with respect to the CG input.
Figure 8: Qualitative comparison of autoregressive long video generation.
B.1 Analysis of Naive DINO Latent Injection
@@
B.2 Impact of Minor Components Pruning Degree ($k^{\prime}$)
@@
B.3 Effect of Spatial Upscaling Factor ($S$)
@@
Figure 9: Qualitative Comparison w. and w./o. Causal Temporal Aggregator
Figure 10: PCA visualization comparing feature maps upsampled via AnyUp ($\times 4$ and $\times 8$) versus naive input upscaling($\times 4$). The red boxes highlight semantic inconsistencies introduced by AnyUp, where high-frequency pixel information leads to false semantic boundaries (e.g., misclassifying the curb). The green boxes demonstrate the failure of AnyUp to recover fine-grained structural details, such as the boundaries of distant street lamps, despite the high upsampling factor.
B.4 Efficacy of the Causal Temporal Aggregator
@@ Appendix C Upsampling using Anyup
While feature upsampling methods such as AnyUp [46] appear to offer a computationally efficient alternative for obtaining high-resolution DINO feature maps, we empirically find that they consistently lead to unstable controllability in video generation, resulting in the suboptimal quantitative metrics and poor visual quality. To investigate this, we visualize the PCA components of feature maps upsampled via AnyUp versus naive input upscaling in Fig. 10.We observe that since AnyUp incorporates high-resolution RGB images as guidance, it inadvertently introduces unnecessary high-frequency pixel information, leading to feature quality degradation. As highlighted by the red boxes, the pixel-level guidance causes AnyUp to misinterpret texture differences (e.g., between the curb and the neighboring ground) as distinct semantic categories, whereas naive upscaling maintains a more coherent semantic representation. Furthermore, as shown in the green boxes, AnyUp fails to recover fine-grained structural details despite an $8\times$ upsampling factor; for instance, the boundaries between distant street lamps and the sky remain indistinct. These visual observations corroborate our experimental findings, confirming that the feature degradation inherent in AnyUp is the primary cause of the generation instability.
@@ Appendix D Robust Long Video Generation using DwD
Appendix E Extra Qualitative Comparisons
Figure 11: Extra Qualitative comparison with state-of-the-art methods. The red boxes highlight inconsistencies with respect to the CG input, while the green boxes point out low-fidelity textures.
Figure 12: Extra Qualitative comparison with state-of-the-art methods. The red boxes highlight inconsistencies with respect to the CG input, while the green boxes point out low-fidelity textures.